Circular No. 6239

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ad IN NGC 2139 R. Evans, Coonabarabran, N.S.W., reports his visual discovery of a

supernova located 25′′ west and 5′′ south of the nucleus of NGC 2139 (α = 6^{h}01^{m}.1, δ = −23^{o}40′, equinox 2000.0); on Sept. 28.8 UT, Evans estimated m_{v} = 14.0–14.5, noting that nothing was visible at this location on Aug. 25 (0.41-m reflector).

S. Benetti, European Southern Observatory (ESO); D. Grupe, Universitäts-Sternwarte, Göttingen, report: “Inspection of a fully-reduced CCD spectrogram (range 390–784 nm, resolution 0.4 nm) obtained on Sept. 29.30 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) at La Silla confirms this object as a type-II supernova around maximum. The spectrum consists of a broad-symmetric (FWHM = 11000 km/s) Hα emission, and P-Cyg profiles of Hδ, Hγ, and Hβ. The lines are superimposed on a blue continuum (T_{bb} ∼ 13 000 K). The expansion velocity deduced from the minima of the H lines is ∼ 9500 km/s. The He I 587.6-nm line is also present with a P-Cyg profile, and it shows a slower expansion velocity (∼ 8800 km/s). The velocity shift (1790 km/s) of a narrow Hα emission component is consistent with the recession velocity of NGC 2139.”

GRO J2058+42 J. E. Grove, Naval Research Laboratory, reports on behalf of the

Compton Gamma Ray Observatory (CGRO) OSSE team: “We have further constrained the position for GRO J2058+42 using the OSSE instrument on the CGRO. Data from a 1-dimensional scan through the source region were combined with the published BATSE error box (IAUC 6238) to yield a new 95-percent confidence region of size ∼ 30′ × 60′, bounded by a quadrilateral with the following vertices (equinox 2000.0): α = 21^{h}00^{m}.5, δ = +41^{o}42′; 20^{h}59^{m}.5, +42^{o}12′; 20^{h}57^{m}.5, +41^{o}21′; 20^{h}58^{m}.3, +40^{o}51′. Follow-up observations in other wavebands to identify the object are encouraged.”

COMET P/1995 S1 (DE VICO) Visual m_{1} estimates: Sept. 22.29 UT, 6.3 (A. Dimai, Cortina d’Ampez-

zo, Italy, 0.50-m reflector); 23.46, 6.4 (W. Dillon, Missouri City, TX, 7×50 binoculars); 24.11, 6.2 (K. Sárneczky, Ráktanya, Hungary, 20×60 binoculars); 25.86, 6: (D.-q. Zhang, China; independent discovery); 26.21, 5.7 (A. Pereira, Cabo Ruivo, Portugal, 9×34 binoculars); 27.17, 5.9 (S. Baroni, Milan, Italy, 20×80 binoculars); 28.43, 5.8 (G. W. Kronk, Troy, IL, 20×80 binoculars); 29.18, 5.4 (J. D. Shanklin, Cambridge, England, naked eye).

1995 September 29 Daniel W. E. Green